OpenStudy (anonymous):

What is the exact value of the expression 5sqrt13+sqrt 6 - 7sqrt 13

OpenStudy (mathstudent55):

\( \sqrt{13} \) and \(\sqrt{6} \) are simple radicals that can't be simplified any further. Treat them as variables. You can combine like variables. \( 5\sqrt{13}\) and \(-7\sqrt{13} \) are like terms. \(\sqrt{6} \) is not a like term to \(\sqrt{13} \).

OpenStudy (anonymous):

Now wat

OpenStudy (mathstudent55):

If you had \( 5x + y - 7x\), you'd be able to combine the terms with x together, but since there is only one term with y, it would remain the same, so \( 5x + y - 7x = -2x + y\) Treat \( \sqrt{13} \) as x, and treat \(\sqrt{6} \) as y.